Presume you have a pump that is increasing the pressure of the liquid passing through it. Assuming the flow is incompressible, what happens to the outlet velocity compared to the inlet when…
a. Diameter decreases
b. Diameter is constant
c. Diameter increases
Answer #1

Solution :

For a fluid flow through a pipe

Area * velocity = constant

A * v = K

  • a. If diameter decrease then area decrease so from above equation A * v = K = A2 * v2 so,
  • v2 = A * v / A2,
  • So from above equation velocity increase.
  • b. If diameter is constant then area remains constant so from above equation A * v = K = A * v2 so,
  • v2 = A * v / A,
  • v2 = v
  • So from above equation velocity remains constant
  • c. If diameter increase then area increase so from above equation A * v = K = A2 * v2 so,
  • v2 = A * v / A2,
  • So from above equation velocity decrease.